2 litres of mercury at 80 deg.Celsius is pushed into a plastic bucket containing 10 litres of mercury at a temperature "T".if the final temperature of the mixture is 30 degree Celsius ,what is the value of "T".(neglect any heat loss to environment)
1
Expert's answer
2011-02-03T07:26:07-0500
We can use the heat balance equation
Cm m1 ΔT1 = Cm m2 ΔT2
ΔT1 = 80 - 30 =50


ΔT2 = 30 - T

m1/m2 = 2/10 = 0.2
Thus

(30-T) = 50*0.2 = 10
T = 30-10 = 20 C.
Answer: T= 20 C.